Session: Feb 2005 (session 1)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Robledo (Cristina Garcia Miro at Tue Mar 1 23:58:22 2005) Minor failures : RFI
Severe RFI on VC03. Weather condition: overcast Station Time delay: -1.6575 microsecs Recorder used: #1 Label(/s): VLBA1224,F23C Calibration at zenith: L-Band System Temp: 21.6 L-Band Y-factor: 11.33 Noise Diodes: 9.7 Ambient Load: 14.312

Shanghai (Tao An at Wed Mar 2 00:30:25 2005) Minor failures :
We lost 4 scans (#0048-0051) due to MK5 problem. error message : ERROR m5 -104 mk5cn: time-out on mk5 device, connection closed. (we also upload the error information upto vlbeer for debugging.) The weather is clear during whole experiment.

Medicina (Andrea Orlati at Wed Mar 2 00:39:54 2005) Minor failures :
Mark5 software hang up after first scan. It did not accept command from both Field System and tstMark5A console. Some scans later the problem step up again. It seems that the reason was due to the disk-pack: one pair of master/slave didn't work properly. After that the run finished without any other problems. Scans 2-13 and 23-35 was lost, included then the first and third ftp scans. Second ftp scan was uploaded to JIVE.

Hartebeesthoek (Sarah Buchner at Wed Mar 2 01:40:20 2005) Success :

Urumqi (Liu Xiang at Wed Mar 2 02:22:30 2005) Minor failures :
No P-cal, it is clear. There is some interferences. The receiver power supply failure maybe at the end of the experiment which caused very low total power spectrum. Disk2file is success, no problem found with using Mark5A normal parsing mode.

Urumqi (Liu Xiang at Wed Mar 2 02:31:50 2005) Minor failures :
We have to stop the experiment at about UT 22:14 because the disk is nearly full. We recorded N05k1, N05c1 and N05l1 in the same disk-pack.

Torun (K. Borkowski at Wed Mar 2 07:07:27 2005) Success :
No known problems.

Onsala (Michael Lindqvist at Wed Mar 2 12:29:30 2005) Minor failures :
Receiver problems. We are using warm HEMTs.

Westerbork (Tony Foley at Thu Mar 3 10:39:52 2005) Minor failures : WIND HIGH
Wb used in phased array mode. Snow at the end of the experiment. This raised the Tsys as the dish was covered. Also one telescope (RT4) has oscillation in the receiver so removed from the tied array

Jodrell_Bank (Alastair Gunn at Mon Mar 7 11:54:52 2005) Success :
The Lovell Telescope was used. No known problems.

Cambridge (Alastair Gunn at Mon Mar 7 11:55:44 2005) Minor failures : RECO
Mk5 unit doesn't seem have recorded scan 2, reason unknown. No other known problems.

Noto (P. Cassaro at Mon Mar 7 15:25:58 2005) Minor failures : POIN
Some pointing problems due to heavy wind. The extraction of the ftp files has failed. The files have been sent after the end of the experiment.

Effelsberg (A. Kraus at Tue Mar 8 12:36:59 2005) Success :